Find the PW of a 10-year bond paying 6% per year nominal rate semiannually that is redeemable at face value if bought by a purchaser to yield 10% per year. The face value of the bond is $1,000. Find out the value of the variables listed below.

  • N = ?
  • r% = ?
  • i% = ?
  • C = ?
  • Z = ?
Answer #1

N = 10 * 2 = 20 periods

r% = 6% / 2 = 3%

i% = 10% / 2 = 5%

C = 3% * 1000 = 30

Z = 30 * (P/A, 5%,20) + 1000 * (P/F, 5%,20)

= 30 * 12.462210 + 1000 * 0.376889

= 750.76